OpenStudy (anonymous):

Find the inverse laplace transform of 1/s^4.

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

ok so laplace transform of t^3 is 6/s^4 so in that fraction we are missing only the 6 so nothing would change in the fraction if we rewrite it as \[\frac{ 1 }{ 6 }\frac{ 6 }{ s^{4} }\] now find inverse laplace of that with pulling out of the inverse laplace the 1/6 since its a constant

OpenStudy (anonymous):

But how did you get the t^3?

OpenStudy (anonymous):

its table laplace transform laplace of t^n is : \[L(t^{n})=\frac{ n! }{ s^{n+1} }\] but here its inverse laplace so you are given the right side so you work backwards

OpenStudy (anonymous):

So you get it from 4-1=3, t^3, right?

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

And how did you get that the laplace transform of t^3 is 6/s^4?

OpenStudy (anonymous):

it is a table laplace transform , i was using the formula i wrote above

OpenStudy (anonymous):

check in you table of laplace transforms you should have laplace transform of t^n

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

ur welcome :)